what is the pulse pressure and mean aterial pressure of a blood pressure of 126/81?
Answer #1

Here, systolic blood pressure,SBP =126 mmHg

Diastolic blood pressure,DBP =81 mmHg

Pulse pressure =SBP - DBP

=(126 - 81) mmHg

=45 mmHg

Mean Arterial pressure,MAP =DBP + P​​​​​​​​​​​ulse pressure/3

=(81 + 45/3)mmHg

=(81+15) mmHg

=96 mmHg.